ebjohnson1:

Michael charges $5 an hour to babysit. Which of the following expresses the amount of money Michael gets paid as a function of how many hours he babysits?

Mercury:

f(x) = 5x f(x) = x + 5 f(x) = x f(x) = x/5 major hint: if he gets 5 hours to babysit, then he makes 5*1 = $5 babysitting one hour; 2*5 = $10 babysitting for 2 hours, etc. so clearly the hourly wage * number of hours is total payment